OpenStudy (anonymous):

Anyone in CA in SC, Algebra 1 B? I need help with Unit 6, Lesson 1.

OpenStudy (anonymous):

Is there a particular question you're stuck on? Or a concept I can try to explain to you? I'm not in your class type, but I could try to explain the concept.

OpenStudy (anonymous):

Organizing Data Using Matrices!

OpenStudy (anonymous):

Umm. Hmm. What type of data are you trying to organize? Give me an example problem please

OpenStudy (anonymous):

Alright, are you working on matrix operations then? For the problem above, when adding matrices we simply add the terms in corresponding locations: \[\left[\begin{matrix}a_{11} & a_{12} \\ a_{21} & a_{22}\end{matrix}\right]+\left[\begin{matrix}b_{11} & b_{12} \\ b_{21} & b_{22}\end{matrix}\right]=\left[\begin{matrix}a_{11}+b_{11} & a_{12}+b_{12} \\ a_{21}+b_{21} & a_{22}+b_{22}\end{matrix}\right]\]

OpenStudy (anonymous):

For the question I got D. Is that correct? I added them. & thanks by the way! :)

OpenStudy (anonymous):

Sorry about the wait, I was engrossed in geometry. Answer D is in fact correct.